# Pure C*-algebras

### v1

@@ Abstract.
We show that every $C^{*}$-algebra that is $(m,n)$-pure in the sense of Winter is already pure, that is, its Cuntz semigroup is almost unperforated and almost divisible. More generally, we show that even weaker comparison and divisibility properties automatically lead to pureness.
As an application to the non-simple Toms-Winter conjecture, we show that every $C^{*}$-algebra with the Global Glimm Property and finite nuclear dimension is pure.
@@
RA and FP were partially supported by Spanish Research State Agency (grant No.  PID2020-113047GB-I00/AEI/10.13039/501100011033), by the Comissionat per Universitats i Recerca de la Generalitat de Catalunya (grant No. 2021-SGR-01015) and by the Spanish State Research Agency through the Severo Ochoa and María de Maeztu Program for Centers and Units of Excellence in R&D (CEX2020-001084-M). HT was partially supported by the Knut and Alice Wallenberg Foundation (KAW 2021.0140).
@@ 1. Introduction
The groundbreaking classification of amenable factors by Connes and Haagerup [Con76, Haa87] in the 1970s and 80s sparked the Elliott program to classify simple, nuclear $C^{*}$-algebras. A first breakthrough in this direction was obtained with the classification in the purely infinite case by Kirchberg and Phillips [Kir95, Phi00] in the 1990s. In the stably finite case, the game-changing examples of Toms [Tom08] showed that additional regularity properties have to be assumed for $K$-theoretic classification, and building on decades of research by numerous mathematicians, in a series of spectacular results [Win14, EGLN15, TWW17, GLN20], the classification of simple, nuclear $C^{*}$-algebras that are $\mathcal{Z}$-stable and satisfy the universal coefficient theorem (UCT) was accomplished in the 2010s. (See also the recent approach [CGS+23] and the survey [Whi23].)
While the UCT holds in all cases of interest (and possibly it is automatic for all nuclear $C^{*}$-algebras), it is known that $\mathcal{Z}$-stability can fail for simple, nuclear $C^{*}$-algebras, and it becomes highly relevant to determine when precisely this regularity property holds. To investigate this problem, Winter [Win12] introduced the notion of (m,n)-pureness for simple $C^{*}$-algebras, with $m$ and $n$ quantifying certain comparison and divisibility properties among operators, and he showed that every simple $C^{*}$-algebra that has locally finite nuclear dimension (a strong form of nuclearity) and that is $(m,n)$-pure for some $m,n$ is $\mathcal{Z}$-stable. Building on work of Rørdam [Rør04], Winter also showed that every simple, $\mathcal{Z}$-stable $C^{*}$-algebra is $(0,0)$-pure (abbreviated to just pure), which means that it satisfies the strongest of the comparison and divisibility properties.
The concept of $(m,n)$-pureness does not depend on the ideal structure of the algebra, hence may be stated for non-simple $C^{*}$-algebras; see Definition 5.1. However, as we show in this paper, all versions of pureness coincide. Our first main result is (see Theorem 5.8):
Theorem A.
If a $C^{*}$-algebra is $(m,n)$-pure for some $m,n\in{\mathbb{N}}$, then it is pure.
This should be contrasted with the well-known examples of simple, amenable $C^{*}$-algebras that do not have 00-comparison ([Vil98]), or that do not have 00-almost divisibility ([DHTW09]). Thus Theorem A can be interpreted as saying that if a $C^{*}$-algebra is not pure, then it necessarily has poor comparison or poor divisibility properties.
Previously, Theorem A has been obtained for separable, simple, unital $C^{*}$-algebras with locally finite nuclear dimension [Win12], and recently for arbitrary simple $C^{*}$-algebras [APRT24, Theorem 10.5]. Our proof of the general non-simple, non-nuclear case is based on the methods developed in [APRT24] in combination with new techniques about what we term functional divisibility (see Section 4 for details). In turn, these are partly inspired by results in [RT17].
Thus a key ingredient in our approach is the Cuntz semigroup and its technical features (see Section 2 for the precise definition). This is an invariant constructed by Cuntz in [Cun78], which is modelled after the Murray-von Neumann semigroup of projections but using positive elements instead. The version we use is what is commonly called the complete Cuntz semigroup, considered in [CEI08], and technically more suitable as it behaves well with respect to, for example, inductive limits. The power of the Cuntz semigroup was demonstrated in the context of the classification program: A. S. Toms produced in [Tom08] two $C^{*}$-algebras that were indistinguishable by the Elliott invariant and a number of reasonable extensions of it, but could be told apart using their Cuntz semigroup. The salient distinguishing feature was strict comparison of positive elements. Another, more recent use of this semigroup is related to the structure of $C^{*}$-algebras of stable rank one, carried out in [APRT22], where powerful Cuntz semigroup techniques yielded solutions to three structural open problems.
The Toms-Winter conjecture
@@
This prominent conjecture [Win18] aims at establishing a robust class of $C^{*}$-algebras covered by the classification theorem referred to above. It predicts that for a separable, simple, unital, non-elementary, nuclear $C^{*}$-algebra, the following properties are equivalent:
@@
    $A$ has finite nuclear dimension.
@@
    $A$ has strict comparison of positive elements.
The nuclear dimension was introduced in [WZ10] as a noncommutative generalization of covering dimension. A $C^{*}$-algebra $A$ is said to be $\mathcal{Z}$-stable if it tensorially absorbs the Jiang-Su algebra $\mathcal{Z}$, that is, if $A\cong\mathcal{Z}\otimes A$. A $C^{*}$-algebra $A$ has strict comparison of positive elements if it has 00-comparison; see Definitions 3.1 and 5.1 for details.
In two groundbreaking papers [Win12, CET+21], it was shown that conditions (1) and (2) of the conjecture are equivalent. This was later generalized to stably projectionless, simple $C^{*}$-algebras in [Tik14, CE20], thereby showing that a separable, simple, non-elementary $C^{*}$-algebra has finite nuclear dimension if and only if it is nuclear and $\mathcal{Z}$-stable.
Further, it is known that (2) implies (3) [Rør04]. The implication (3$\Rightarrow$(2) has been verified under certain additional assumptions [Sat12, KR14, TWW15, Thi20], but it remains open in general.
The Toms-Winter conjecture is also closely related to pureness, which we view as a property lying between (2) and (3). Indeed, as mentioned above, $\mathcal{Z}$-stability implies pureness, which in turn implies strict comparison of positive elements; see Proposition 5.2. Further, among separable, simple, unital, non-elementary $C^{*}$-algebras, pureness implies $\mathcal{Z}$-stability under the additional assumption of locally finite nuclear dimension [Win12], and strict comparison of positive elements implies pureness under the additional assumption of a positive solution to the Rank Problem. Such a positive solution has been obtained for $C^{*}$-algebras of stable rank one [Thi20, APRT22].
The non-simple Toms-Winter conjecture
The regularity properties of the Toms-Winter conjecture are also relevant for non-simple $C^{*}$-algebras, and an investigation of their interplay was initiated in [RT17]. While the assumption of unitality in the original formulation of the conjecture could be dispensed with, the condition to be non-elementary is crucial. Indeed, elementary $C^{*}$-algebras have nuclear dimension zero, but they are neither $\mathcal{Z}$-stable nor pure.
There are two natural generalizations of non-elementariness to the non-simple setting: Nowhere scatteredness, and the Global Glimm property. Here, a $C^{*}$-algebra is nowhere scattered if none of its quotients has a nonzero, elementary ideal, or equivalently if every hereditary sub-$C^{*}$-algebra is generated by nilpotent elements [TV24]. A $C^{*}$-algebra has the Global Glimm Property if every hereditary sub-$C^{*}$-algebra contains an approximately full, nilpotent element [TV23]. Clearly, every $C^{*}$-algebra with the Global Glimm Property is nowhere scattered, and the Global Glimm Problem asks whether the converse holds. We ask:
Question B.
@@
The implications (2$\Rightarrow$(3a)$\Rightarrow$(3b) hold in general; see Proposition 5.2. We also note that pureness implies the Global Glimm Property; see [TV23, Theorem 3.6]. Therefore, Question B contains in particular the question of whether the Global Glimm Problem has a positive solution for $C^{*}$-algebras with finite nuclear dimension. As for simple $C^{*}$-algebras, the question whether strict comparison of positive elements implies pureness is closely related to the Rank Problem. This was solved for $C^{*}$-algebras of real rank zero or stable rank one ([ER06] and [APRT22]), which shows that (3b) implies (3a) in these cases.
In [RT17, Theorems 7.10, 7.15] it is shown that (3a) implies (2) under the additional assumption that the $C^{*}$-algebra $A$ has locally finite nuclear dimension, that no quotient of $A$ has a simple, purely infinite ideal, and that the primitive ideal space of $A$ has a basis of compact-open subsets or is Hausdorff.
The following result (which is Theorem 6.4) provides a further partial answer to Question B.
Theorem C.
@@ Acknowledgements
The authors thank Leonel Robert for valuable feedback on an earlier version of this paper.
@@ 2. Preliminaries
The techniques used in the paper mainly concern the structure of the Cuntz semigroup $\operatorname{Cu}(A)$ of a $C^{*}$-algebra $A$, whose definition and properties we describe below.
@@
It is proved in [CEI08] that $\operatorname{Cu}(A)$ satisfies the following properties defined for a general positively ordered monoid $S$ (and usually referred to as axioms):
@@
    If $(x_{n})_{n}$ is an increasing sequence in $S$, then $\sup_{n}x_{n}$ exists.
@@
    For any $x\in S$ there exists a sequence $(x_{n})_{n}$ such that $x_{n}\ll x_{n+1}$ for all $n$ and $x=\sup_{n}x_{n}$. (We say that $(x_{n})_{n}$ is a $\ll$-increasing sequence.)
@@
    If $(x_{n})_{n}$ and $(y_{n})_{n}$ are increasing sequences in $S$, then $\sup_{n}(x_{n}+y_{n})=\sup_{n}x_{n}+\sup_{n}y_{n}$.
The relation $\ll$ in these axioms is defined as follows: $x\ll y$ if for every increasing sequence $(y_{n})_{n}$ satisfying $y\leq\sup_{n}y_{n}$ there exists $n_{0}\in{\mathbb{N}}$ such that $x\leq y_{n_{0}}$. The relation $\ll$ is called the way-below relation, or compact containment relation, and one says that ‘$x$ is way-below $y$’ if $x\ll y$. An element $u\in S$ such that $u\ll u$ is termed compact.
A positively ordered monoid satisfying axioms (O1)-(O4) is called a $\mathrm{Cu}$-semigroup and belongs to a category of semigroups that has thoroughly been studied (see, for example, [APT18], [APT20a], [APT20b] [APT20c], [TV22], [APRT22], among others; see also the recent survey [GP23]). Yet, when dealing with semigroups obtained as Cuntz semigroups of $C^{*}$-algebras, or further restricting to particular families of $C^{*}$-algebras, additional axioms and properties are satisfied.
The first axiom that is added after (O4) is the following:
@@
This axiom is usually referred to as the axiom of almost algebraic order. The first part of the statement above was proved in [RW10, Lemma 7.1], and the second part can be found in [APT18, Proposition 4.6].
@@
Recall that a $\mathrm{Cu}$-semigroup $S$ is said to be countably based if all elements can be written as suprema of elements from a fixed countable subset. It is known that $\operatorname{Cu}(A)$ is countably based whenever $A$ is separable (for a proof, see [APS11, Lemma 1.3]).
Theorem 2.2 (cf. [APRT21, Theorems 2.4, 2.5]).
@@
Given $x\in S$, we obtain a map $\widehat{x}\colon F(S)\to[0,\infty]$ defined by $\widehat{x}(\lambda)=\lambda(x)$, for $\lambda\in F(S)$. $\widehat{x}$ is linear and lower semicontiuous, and if $S$ satisfies (O5) and (O6), then $\textrm{Lsc}(F(S))$, the set of all such maps, is an algebraically ordered, lattice ordered, compact cone.
Two main questions arise in this setting. To determine the range of the so-called rank map $x\mapsto\widehat{x}$ (referred to as the Rank Problem); and determine to what extent the order in $S$ can be inferred by the order in $\textrm{Lsc}(F(S))$. Note that $x\mapsto\widehat{x}$ is additive and preserves sequential suprema; it does not preserve the relation of compact containment, and is certainly not an order embedding. However, the result below will be useful.
Given elements $x$ and $y$ in a partially ordered semigroup, one writes $x<_{s}y$ if there exists $k\in{\mathbb{N}}$ such that $(k+1)x\leq ky$. Item (ii) in the following lemma is a consequence of [Rob13, Proposition 2.2.2] or [APT18, Theorem 5.2.19].
@@
    If $\widehat{x}\leq\gamma\widehat{y}$ for some $\gamma\in(0,1)$, then $x^{\prime}<_{s}y$ for all $x^{\prime}\ll x$.

### v3

@@ Abstract.
We demonstrate that pure $C^{*}$-algebras form a robust class by proving that pureness follows from very weak comparison and divisibility properties. Using this, we show that every simple, non-elementary $C^{*}$-algebra with a unique quasitrace and with very mild comparison is pure, and, as a result, has strict comparison. Furthermore, sufficiently non-commutative $C^{*}$-algebras of stable rank one and with weak comparison are likewise pure.
We also show that adequately non-elementary $C^{*}$-algebras with finite nuclear dimension are pure, which leads to the verification of the non-simple Toms-Winter conjecture for a large class of $C^{*}$-algebras.
@@
RA, FP and EV were partially supported by the Spanish Research State Agency (grant No.  PID2020-113047GB-I00/AEI/10.13039/501100011033) and by the Comissionat per Universitats i Recerca de la Generalitat de Catalunya (grant No. 2021-SGR-01015). RA and FP were also partially supported by the Spanish State Research Agency through the Severo Ochoa and María de Maeztu Program for Centers and Units of Excellence in R&D (CEX2020-001084-M). HT and EV were partially supported by the Knut and Alice Wallenberg Foundation (KAW 2021.0140). EV was also supported by the Fields Institute for Research in Mathematical Sciences.
@@ 1. Introduction
Pureness is a regularity property for $C^{*}$-algebras introduced by Winter in his seminal investigation into $\mathcal{Z}$-stability and finite nuclear dimension of simple, nuclear $C^{*}$-algebras [Win12]. Interest in these concepts was sparked by Toms’ groundbreaking examples [Tom08], where he constructed two simple, nuclear $C^{*}$-algebras with the same $K$-theoretic and tracial data —one pure, $\mathcal{Z}$-stable (i.e., tensorially absorbing the Jiang-Su algebra $\mathcal{Z}$), and of finite nuclear dimension, and the other not. This led to a revision of the Elliott program, aiming at classifying separable, simple, nuclear $C^{*}$-algebras. Building on decades of research by numerous mathematicians, this was ultimately achieved in a series of remarkable results [Win14, TWW17, GLN20]. (See also the recent approach [CGS+23] and the survey [Whi23].) The classification theorem states that separable, simple, nuclear $C^{*}$-algebras that are $\mathcal{Z}$-stable and satisfy the universal coefficient theorem (UCT) can be distinguished up to isomorphism by their Elliott invariant, which essentially consists of topological K-Theory and the tracial simplex.
As defined by Winter, a $C^{*}$-algebra is pure if it has strict comparison and is almost divisible. These are $C^{*}$-algebraic analogs of two fundamental properties of projections in $\mathrm{II}_{1}$ factors discovered by Murray and von Neumann: first, the comparison of projections in a $\mathrm{II}_{1}$ factor is determined by its unique trace $\tau$. Second, every projection $p$ is divisible in the sense that for every $n\geq 2$ there exists a projection $q$ with $\tau(q)=\tfrac{1}{n}\tau(p)$. Strict comparison for $C^{*}$-algebras was first considered by Blackadar [Bla88], and is currently a well-established concept in the structure theory of $C^{*}$-algebras. This property is difficult to prove and its validty has far-reaching implications (see, for example, its role in the Toms-Winter conjecture below 1.2). In fact, while comparison and divisibility hold automatically in every $\mathrm{II}_{1}$ factor, both strict comparison and almost divisibility can fail in simple $C^{*}$-algebras [Vil98, Vil99, DHTW09].
With these parallels in mind, one can think of finite, simple, pure $C^{*}$-algebras as $C^{*}$-algebraic analogs of $\mathrm{II}_{1}$ factors. The concept also makes sense for infinite (that is, not finite) $C^{*}$-algebras, and it turns out that a simple $C^{*}$-algebra is purely infinite if and only if it is infinite and pure, and one can think of such algebras as the $C^{*}$-algebraic analogs of $\mathrm{III}$ factors. More generally, nonsimple pure $C^{*}$-algebras may be considered as analogs of von Neumann algebras with trivial type $\mathrm{I}$ summand.
Nowadays, the importance of pureness goes beyond its original intent. For separable $C^{*}$-algebras, pureness of the central sequence algebra [KR14] is equivalent to $\mathcal{Z}$-stability of the underlying $C^{*}$-algebra ([PTV24]); for non-nuclear $C^{*}$-algebras —such as those coming from non-amenable groups or certain infinite reduced free products— pureness becomes an instrument weaker than $\mathcal{Z}$-stability for establishing strict comparison; for a $C^{*}$-algebra with mild divisibility (resp. comparison) properties, showing that it is not pure implies that is has very poor comparison (resp. divisibility) in the precise sense of Theorem C below. Due to these and other applications, the study of pureness has produced many important open questions: Is there a non-simple version of the Toms-Winter conjecture (1.3)? Do all $\mathrm{C}^{*}$-simple groups have strict comparison? (Equivalently, are they all pure? It is known that many are, as shown in [AGKP24], but in general the question is open.) Are there real rank zero versions of Rørdam’s example [Rør03]? (Conversely, are all non-elementary, simple, real rank zero $C^{*}$-algebras pure?)
It is thus of utmost importance to understand which $C^{*}$-algebras are pure. In this paper, we provide a new approach to the study of pureness by analyzing two new (and considerably weak) conceptual properties which we term controlled comparison and functional divisibility. The study of these notions —discussed in 1.1 below— allows us, among other things, to prove robustness results for the class of pure $C^{*}$-algebras. One of the main applications of this new approach is the uncovering of new families where pureness can be verified. (The notions of nowhere scatteredness and the Global Glimm property mentioned below can be thought of as suitable generalizations of non-elementariness to the non-simple case; see the discussion before; see the discussion before Question D.)
Theorem A (5.12, 5.13).
Let $A$ be a $C^{*}$-algebra with controlled comparison. Moreover, assume that
  1. (i)
    $A$ is nowhere scattered with real rank zero or stable rank one, or
  2. (ii)
    $A$ is simple, unital, non-elementary, with a unique quasitracial state.
Then $A$ is pure. In particular, $A$ has strict comparison.
Theorem B (6.5).
Theorem B naturally fits in the context of the non-simple Toms-Winter conjecture (1.3), while Theorem A can be applied to large families of $C^{*}$-algebras: The conditions in (ii) are satisfied by any exact $\mathrm{C}^{*}$-simple group ([BKKO17]) and by Villadsen algebras of the second type [Vil99], while (i) is known to hold for all Villadsen algebras of the first type [Vil98]. Thus, exact $\mathrm{C}^{*}$-simple groups and Villadsen algebras have strict comparison if and only if they have controlled comparison, if and only if they are pure.
As previously stated, it is an important open problem to determine if all non-elementary simple real rank zero $C^{*}$-algebras have strict comparison (or if they are $\mathcal{Z}$-stable under the assumption of nuclearity and separability). Part (i) of Theorem A also applies to this class of algebras, reducing the problem to proving controlled comparison.
1.1. Controlled comparison and functional divisibility
Together with his notion of pureness, Winter introduced in [Win12] the quantified versions of $(m,n)$-pureness, where $m$ and $n$ are non-negative integers measuring how big the failure of comparison ($m$) and divisibility ($n$) is. Surprisingly, it follows from Winter’s work that for simple $C^{*}$-algebras with locally finite nuclear dimension there is a dimension reduction phenomenon: $(m,n)$-pureness automatically implies $(0,0)$-pureness, that is, pureness. This was recently extended to the simple (possibly non-nuclear) setting in [APRT24, Theorem 10.5]. This dimension reduction phenomenon is one of the main techniques for establishing pureness or strict comparison outside the $\mathcal{Z}$-stable case, and a consequence of our Theorem C is that it holds for all $C^{*}$-algebras.
The novelty of this paper lies in the study of two unquantified notions which we call controlled comparison and functional divisibility. Each of these notions is weaker than any of their quantified counterparts, thus making them easier to check in concrete examples.
Inspired by the study initiated in [RT17], we show that there is a beautiful interplay between these functional notions and relative central sequences. This ultimately allows us to bypass any ideal-lattice obstruction and obtain our main tool for verifying pureness:
Theorem C (5.7).
Given a $C^{*}$-algebra $A$, the following statements are equivalent:
  1. (i)
    $A$ is pure, that is, it has strict comparison and is almost divisible.
  2. (ii)
    $A$ has controlled comparison and is functionally divisible.
Theorem C implies, in particular, that the dimension reduction phenomenon alluded to above holds in full generality: If a $C^{*}$-algebra is $(m,n)$-pure for some $m$ and $n$, then it is pure.
The present result provides a general statement that can be applied to non-simple, non-nuclear, $C^{*}$-algebras such as, for example, central sequence algebras. Even in the simple case, Theorem C represents a substantial improvement of [APRT24, Theorem 10.5]: Given a $C^{*}$-algebra $A$ satisfying condition (i) or (ii) in Theorem A, it is not known if $A$ satisfies the divisibility property in $(m,n)$-pureness (known as $n$-almost divisibility) for any $n$. However, we show in Propositions 4.10 and 4.11 that any $C^{*}$-algebra satisfying condition (i) or (ii) in Theorem A is automatically functionally divisibile.
1.2. The Toms-Winter conjecture
@@
This conjecture [Win18] aims at establishing a class of $C^{*}$-algebras covered by the classification theorem mentioned at the beginning of the introduction. Explicitly, it predicts that for a separable, simple, unital, non-elementary, nuclear $C^{*}$-algebra, the following properties are equivalent:
@@
    The $C^{*}$-algebra has finite nuclear dimension, a noncommutative generalization of finite covering dimension introduced in [WZ10].
@@
    The $C^{*}$-algebra has strict comparison of positive elements (see Section 3).
In two groundbreaking papers [Win12, CET+21], it was shown that conditions (1) and (2) are equivalent. Further, it is known that (2) implies (3) [Rør04]. The implication (3$\Rightarrow$(2) has been verified under certain additional assumptions [Sat12, KR14, TWW15, Thi20], but remains open in general.
The Toms-Winter conjecture is closely related to pureness, which we view as a property lying between (2) and (3). Indeed, $\mathcal{Z}$-stability implies pureness, which in turn implies strict comparison of positive elements; see Proposition 5.2. Further, strict comparison of positive elements implies pureness under the additional assumption of a positive solution to the Rank Problem; see Section 2. Such a positive solution has been obtained for $C^{*}$-algebras of real rank zero [ER06] and of stable rank one [Thi20, APRT22].
1.3. The non-simple Toms-Winter conjecture
The regularity properties of the Toms-Winter conjecture are also relevant for non-simple $C^{*}$-algebras, and an investigation of their interplay was initiated by Robert and Tikuisis [RT17]. While the assumption of unitality in the original formulation of the Toms-Winter conjecture could be dispensed with, the condition of being non-elementary is crucial. Indeed, elementary $C^{*}$-algebras have nuclear dimension zero, but they are neither $\mathcal{Z}$-stable nor pure.
The natural generalization of non-elementariness to the non-simple setting is the notion of nowhere scateredness: a $C^{*}$-algebra is nowhere scattered if none of its quotients has a nonzero, elementary ideal, or equivalently if every hereditary sub-$C^{*}$-algebra is generated by nilpotent elements [TV24a]. This notion is tightly related to the Global Glimm Property: a $C^{*}$-algebra has the Global Glimm Property if every hereditary sub-$C^{*}$-algebra contains an approximately full, nilpotent element [TV23]. Clearly, every $C^{*}$-algebra with the Global Glimm Property is nowhere scattered. The converse remains open, and is known as the Global Glimm Problem. In light of these considerations, the following question is natural:
Question D.
@@ 1.3. The non-simple Toms-Winter conjecture
The implications (2$\Rightarrow$(3a)$\Rightarrow$(3b) hold in general, and pureness always implies the Global Glimm Property; see Proposition 5.2. Therefore, Question D contains in particular the Global Glimm Problem for $C^{*}$-algebras with finite nuclear dimension.
Partial results on the implication (3a)$\Rightarrow$(2) were established in [RT17, Theorems 7.10, 7.15], while the implication (2$\Rightarrow$(1) has been shown to hold in a variety of cases; see, for example, [BL24, BGSW22, ENST20]. As for simple $C^{*}$-algebras, whether (3b) implies (3a) is related to the Rank Problem.
Combining Theorem B with some of the above mentioned results, we obtain an answer to Question D for a relevant class of $C^{*}$-algebras, thereby verifying the non-simple Toms-Winter conjecture in this setting.
Theorem E (6.6).
Let $A$ be a separable, locally subhomogeneous $C^{*}$-algebra that has stable rank one, topological dimension zero, and the Global Glimm Property. Then, the answer to Question D is positive.
Conventions
Given a $C^{*}$-algebra, we let $A_{+}$ to denote its positive elements. We let ${\mathbb{N}}$ denote the set of natural numbers, including 00. Further, ${\mathcal{K}}$ denotes the $C^{*}$-algebra of compact operators on a separable, infinite-dimensional Hilbert space.
@@ Acknowledgements
The authors thank Leonel Robert, Stuart White and Wilhelm Winter for valuable feedback on earlier versions of this paper.
@@ 2. Preliminaries
The techniques used in this paper mainly concern the structure of the Cuntz semigroup of a $C^{*}$-algebra whose definition and properties we describe below.
@@
It is proved in [CEI08] that $\operatorname{Cu}(A)$ satisfies the following properties (usually referred to as axioms):
@@
    If $(x_{n})_{n}$ is an increasing sequence in $\operatorname{Cu}(A)$, then $\sup_{n}x_{n}$ exists.
@@
    For any $x\in\operatorname{Cu}(A)$ there exists a sequence $(x_{n})_{n}$ such that $x_{n}\ll x_{n+1}$ for all $n$ and $x=\sup_{n}x_{n}$. (We say that $(x_{n})_{n}$ is a $\ll$-increasing sequence.)
@@
    If $(x_{n})_{n}$ and $(y_{n})_{n}$ are increasing sequences in $\operatorname{Cu}(A)$, then $\sup_{n}(x_{n}+y_{n})=\sup_{n}x_{n}+\sup_{n}y_{n}$.
The relation $\ll$ in these axioms is defined as follows: $x\ll y$ if for every increasing sequence $(y_{n})_{n}$ satisfying $y\leq\sup_{n}y_{n}$ there exists $n_{0}\in{\mathbb{N}}$ such that $x\leq y_{n_{0}}$. The relation $\ll$ is called the way-below relation, or the compact containment relation, and one says that ‘$x$ is way-below $y$’ (or that ‘$x$ is compactly contained in $y$’) if $x\ll y$. An element $u\in S$ such that $u\ll u$ is termed compact.
A positively ordered monoid satisfying axioms (O1)-(O4) is called a $\mathrm{Cu}$-semigroup and belongs to a category of semigroups that has thoroughly been studied (see, for example, [APT18, APT20c, APT20a, APT20b], among others; see also the recent survey [GP24]). Yet, when dealing with semigroups obtained as Cuntz semigroups of $C^{*}$-algebras, additional axioms and properties are satisfied:
@@
This axiom is usually referred to as the axiom of almost algebraic order. For the Cuntz semigroup of an arbitrary $C^{*}$-algebra, the first part of the statement above was proved in [RW10, Lemma 7.1], and the second part can be found in [APT18, Proposition 4.6].
@@
Recall that a $\mathrm{Cu}$-semigroup $S$ is said to be countably based if all elements can be written as suprema of elements from a fixed countable subset. It is known that $\operatorname{Cu}(A)$ is countably based whenever $A$ is separable; see [APS11, Lemma 1.3].
Theorem 2.2 ([APRT21, Theorems 2.4, 2.5]).
@@
Given $x\in S$, we obtain a map $\widehat{x}\colon F(S)\to[0,\infty]$ defined by $\widehat{x}(\lambda)=\lambda(x)$, for $\lambda\in F(S)$. The map $\widehat{x}$ is linear and lower semicontinuous. Denote by $\operatorname{Lsc}(F(S))$ the set of all linear, lower semicontinuous maps on $F(S)$ with values in $[0,\infty]$.
Two main questions arise in this setting. The first concerns the extent to which the ordering in $A$ can be inferred from the ordering in $\textrm{Lsc}(F(S))$. Note that whilst $x\mapsto\widehat{x}$ is an additive map that preserves the order and suprema of increasing sequences, it typically does not preserve the way-below relation and, moreover, is generally not an order embedding. Still, the result below will be useful.
Given elements $x$ and $y$ in a partially ordered semigroup, one writes $x<_{s}y$ if there exists $k\in{\mathbb{N}}$ such that $(k+1)x\leq ky$.
@@
    If $\widehat{x}\leq\gamma\widehat{y}$ for some $\gamma\in(0,1)$, then $x^{\prime}<_{s}y$ for all $x^{\prime}\in S$ with $x^{\prime}\ll x$.
The second question that arises consists of determining the range of the map $x\mapsto\widehat{x}$. This is referred to as the Rank Problem and was explored in detail in [Thi20] and [APRT22]. For a $\mathrm{Cu}$-semigroup $S$, let us denote by $\textrm{L}(F(S))$ the smallest subsemigroup of $\operatorname{Lsc}(F(S))$ closed under suprema of increasing sequences and containing all elements of the form $\tfrac{1}{n}\widehat{x}$ (see [Rob13, Section 3]). Recall that an element $x\in S$ is termed soft if for every $x^{\prime}\in S$ with $x^{\prime}\ll x$ there exists $t\in S$ such that $x^{\prime}+t\ll x$ and $x^{\prime}\ll\infty t$ (see [TV24b, Definition 4.3, Remark 4.4]; for $C^{*}$-algebras of stable rank one, this agrees with [APT18, Definition 5.3.1]).
The following theorem summarizes the realization results from [APRT22], which will be required in Section 4. The existence of the map $\alpha$ is shown in [APRT22, Theorem 7.2], and the properties of $\alpha$ are proved in [APRT22, Proposition 7.4, Theorem 7.13, Proposition 8.2, Theorem 8.4, Corollary 8.5].
Theorem 2.4 ([APRT22]).
Let $A$ be a separable, nowhere scattered $C^{*}$-algebra with stable rank one. Then there is a natural, additive, order-preserving map $\alpha\colon\textrm{L}(F(\operatorname{Cu}(A)))\to\operatorname{Cu}(A)$ that preserves suprema of increasing sequences and such that $f=\widehat{\alpha(f)}$ for every $f\in\textrm{L}(F(\operatorname{Cu}(A)))$.
The image of $\alpha$ consists of soft elements, and we have
$\alpha(f+\widehat{x})=\alpha(f)+x$
whenever $x\in\operatorname{Cu}(A)$ and $f\in\textrm{L}(F(\operatorname{Cu}(A)))$ satisfy $\widehat{x}\leq\infty f$.